# Adjoint log canonical foliated singularities on surfaces

### v1

@@ 4.4.2. Case II: $E$ is a $K^{=0}$-graph
If $E$ consists of a single component $C$, then $KC=0$ and $d(C)=0$. Thus $C$ falls into case (1-a) of Lemma 4.2 or Lemma 4.3. This is case (1).
@@
By Lemma 4.2 and Lemma 4.3, each $C_{i}$ must belong to one of the following types:
@@
A direct analysis using the separatrix theorem shows that cases (i) and (iii) cannot occur. Therefore, each $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve, and $E$ is a cycle of such curves. This is case (2). ∎
4.4.3. The presence of special $K^{>0}$-chains
We now assume that $E$ does not belong to Case I or Case II. Then $E$ necessarily contains at least one special $K^{>0}$-chain.
@@ 4.4.3. The presence of special $K^{>0}$-chains
Suppose that there exist two special $K^{>0}$-chains $F_{1}$ and $F_{2}$ with distinct first curves, such that they share a common component or satisfy $F_{1}F_{2}>0$. Then $E=C_{1}+\cdots+C_{r}$ is a chain belonging to one of the following cases:
@@
By Proposition 4.5, a simple combinatorial analysis yields the desired cases, and the separatrix theorem excludes all others. ∎
To separate the remaining cases, we impose the following technical assumption.
    $E$ does not belong to Case I or Case II, and any two special $K^{>0}$-chains with distinct first curves are pairwise disjoint.
Under assumption ()(*), maximal special $K^{>0}$-chains can be defined. We introduce the following notation.
  • Let $\{\Theta_{1},\dots,\Theta_{s}\}$ denote the set of all maximal special $K^{>0}$-chains of type (A) contained in $E$.
  • Let $\{\Xi_{1},\dots,\Xi_{t}\}$ denote the set of all maximal special $K^{>0}$-chains of type (B), (C), or (D) contained in $E$.
Define
@@
Under assumption ()(*), let $C$ be an irreducible non-$\mathcal{F}$-invariant curve contained in $E$, and assume that $C$ is not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
@@
    $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$, and $E$ is an $\mathcal{F}$-star graph (cf. Definition 2.22) of type $(l;m_{1},\dots,m_{l})$ centered at $C$, where $(l;m_{1},\dots,m_{l})$ is one of the following:
@@
    (In cases (b) and (c), we have $K_{1}C=0$.)
@@
    In this case, $K_{1}C=0$.
@@
In this case we have $d(C)\geq 1$. If $FC>0$ for some special $K^{>0}$-chain $F$, then $F$ must be of type (C) by Proposition 4.5 and Corollary 4.7.
Without loss of generality, we assume that $C$ meets transversely the chains $\Xi_{1},\dots,\Xi_{l}$. Let $m_{j}$ denote the absolute value of the determinant of the intersection matrix of $\Xi_{j}$. Since each $\Xi_{j}$ is of type (C), for $j\leq l$, we have $m_{j}\geq 2$. By (4.4), we obtain
@@
$\geq{\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\epsilon(d_{2}-l)+\frac{\epsilon l}{2}$
$={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\frac{\epsilon(d_{2}-l)}{2}+\frac{\epsilon d_{2}}{2},$
@@
Note that if ${\rm tang}(\mathcal{F},C)\geq 1$ or $p_{a}(C)\geq 1$ or $d_{1}\geq 1$, then $K_{1}C>0$, where we use $d=d_{1}+d_{2}\geq 1$ and $\epsilon\in(0,1/3)$. Hence we may assume
${\rm tang}(\mathcal{F},C)=p_{a}(C)=d_{1}(C)=0.$
In this case,
$K_{1}C=\epsilon(d_{2}-l-2)+\sum_{j=1}^{l}\epsilon\!\left(1-\frac{1}{m_{j}}\right)\geq\epsilon(d_{2}-l-2)+\frac{\epsilon l}{2}.$
  • If $d_{2}-l\geq 3$, or $d_{2}-l=2$ with $l\geq 1$, or $d_{2}-l=1$ with $l\geq 3$, or $d_{2}-l=0$ with $l\geq 5$, then $K_{1}C>0$.
  • If $d_{2}-l=2$ with $l=0$, then $K_{1}C=0$. Here $C$ is disjoint from all special $K^{>0}$-chains and $d=d_{2}=2$. This is case (3).
  • If $d_{2}-l=1$ with $l=2$, then
@@
    and equality holds iff $(m_{1},m_{2})=(2,2)$. This is case (2).
  • If $d_{2}-l=1$ with $l=1$, then $d=d_{2}=2$ and $KC=0$. In this situation $\Xi_{1}+C$ forms a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $d_{2}-l=1$ with $l=0$, or $d_{2}-l=0$ with $l=1$, then $d=1$ and $KC<0$. Thus $C$ would be the first curve of a special $K^{>0}$-chain, contradicting our assumption.
  • If $d_{2}-l=0$ with $l=0$, then $d=d_{2}=0$, contradicting our assumption.
  • If $d_{2}-l=0$ with $l=2$, then $d=d_{2}=2$ and $KC=0$. In this situation $\Xi_{1}+C$ forms a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $d_{2}-l=0$ with $l\in\{3,4\}$, then
    $K_{1}C=-2\epsilon+\sum_{i=1}^{l}\epsilon(1-\frac{1}{m_{j}})=\epsilon\cdot\left(-2+l-\sum_{i=1}^{l}\frac{1}{m_{j}}\right)$
@@
    In this case $E=C+\Xi_{1}+\cdots+\Xi_{l}$ is an $\mathcal{F}$-star graph, giving case (1).
@@
    In this case, $K_{1}C=0$.
@@
In particular, cases (2-b), (2-c), (3-b) and (3-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
@@
Without loss of generality, we assume $C$ meets transversely with $\Theta_{1},\cdots,\Theta_{k}$ at $p_{1},\cdots,p_{k}$ respectively, and $C$ meets transversely with $\Xi_{1},\cdots,\Xi_{l}$ at $q_{1},\cdots,q_{l}$ respectively. Let
$V:=\sum_{i=1}^{k}M(K,\Theta_{i})+\sum_{j=1}^{l}M(K,\Xi_{i}).$
@@
(4.6) $VC=\sum_{i=1}^{k}\frac{1+\epsilon}{n_{i}}+\sum_{j=1}^{l}\frac{\epsilon}{m_{j}}$
where $n_{i}$ (resp. $m_{i}$) denotes the absolute value of the determinant of the intersection matrix of $\Theta_{i}$ (resp. $\Xi_{j}$).
By (4.3),
@@
$\qquad+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}}),$
@@
If $p_{a}(C)\geq 1$, then $K_{1}C\leq 0$ implies $k=l=d=0$, which is a contradiction with the assumption ()(*).
Next we assume $p_{a}(C)=0$. Then the dual graph of $C+\sum_{i=1}^{k}\Theta_{i}+\sum_{j=1}^{l}\Xi_{j}$ is a tree and
$K_{1}C$ $=(1+\epsilon)(d-k-l-2)+l+{\rm Z}(\mathcal{F},C)-d_{2}$
$\qquad+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}})$
$\geq(1+\epsilon)(d-k-l-2)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\frac{k(1+\epsilon)}{2}$
Suppose $d-k-l\geq 3$. Then $K_{1}C>0$.
@@
Suppose dkl=2d-k-l=2. Then $K_{1}C\leq 0$ implies $k=l=0$, $d=2$, and ${\rm Z}(\mathcal{F},C)=d_{2}\in\{1,2\}$. Thus, $C$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$ or a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=2$, and it is disjoint from all special $K^{>0}$-chains. This is case (4) or (5). In this case, $K_{1}C=KC=0$.
Suppose dkl=1d-k-l=1. Then
$K_{1}C$ $=-(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}})$
$\geq-(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\tfrac{k}{2}(1+\epsilon).$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}\geq 2$, then
    $K_{1}C\geq-(1+\epsilon)+2=1-\epsilon>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k\geq 1$, then
    $K_{1}C\geq-(1+\epsilon)+1+\frac{1}{2}(1+\epsilon)=\frac{1}{2}(1-\epsilon)>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k=0$, then $d=l+1\in\{1,2\}$.
    • If $l=0$, then $d=1$ and ${\rm Z}(\mathcal{F},C)-d_{2}=1$. Thus, $KC=-\epsilon$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $l=1$, then $d=2$ and ${\rm Z}(\mathcal{F},C)-d_{2}=0$. Thus, $KC=0$. In this case, $\Xi_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $l={\rm Z}(\mathcal{F},C)-d_{2}=0$, then
    $K_{1}C=-(1+\epsilon)+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})\geq(1+\epsilon)(\frac{k}{2}-1).$
    Thus, $K_{1}C\leq 0$ implies $k\in\{0,1,2\}$.
    • If $k=0$, then $d=1$ and $KC=-(1+\epsilon)$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $k=1$, then $d=2$ and $KC=0$. In this case, $\Theta_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Theta_{1}$.
    • If $k=2$, then $d=3$ and
      $K_{1}C=(1+\epsilon)(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}).$
      Hence, $K_{1}C\leq 0$ implies $(n_{1},n_{2})=(2,2)$. Thus, $C$ meets transversely two (1)(-1)-$\mathcal{F}$-curves $\Gamma_{1}$ and $\Gamma_{2}$ with
      $d(\Gamma_{1})=d(\Gamma_{2})=1,\qquad\Gamma_{1}^{2}=\Gamma_{2}^{2}=-2.$
      By the separatrix theorem, we see $d_{2}(C)={\rm Z}(\mathcal{F},C)>2$. So $d_{2}(C)=d(C)=3$. This is case (3). In this case, $K_{1}C=0$.
Suppose dkl=0d-k-l=0. Then
$K_{1}C$ $=-2(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\sum_{i=1}^{k}(1+\epsilon)(1-\frac{1}{n_{i}})+\sum_{j=1}^{l}\epsilon(1-\frac{1}{m_{j}})$
$\geq-2(1+\epsilon)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\tfrac{k}{2}(1+\epsilon).$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}\geq 3$, then
    $K_{1}C\geq-2(1+\epsilon)+3=1-2\epsilon>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=2$ and $k\geq 1$, then
    $K_{1}C\geq-2(1+\epsilon)+2+\frac{1}{2}(1+\epsilon)=\frac{1}{2}(1-3\epsilon)>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=2$ and $k=0$, then $d=l\in\{1,2\}$.
    • If $l=1$, then $d=1$ and ${\rm Z}(\mathcal{F},C)-d_{2}=1$ and so $KC=-\epsilon$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $l=2$, then $d=2$ and ${\rm Z}(\mathcal{F},C)-d_{2}=0$ and so $KC=0$. In this case, $\Xi_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k\geq 3$, then
    $K_{1}C\geq-2(1+\epsilon)+1+\frac{3}{2}(1+\epsilon)=\frac{1}{2}(1-\epsilon)>0.$
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k=2$, then $d=l+2\in\{2,3\}$ and
    $K_{1}C\geq-2(1+\epsilon)+1+\sum_{i=1}^{2}(1+\epsilon)(1-\frac{1}{n_{i}})=1-(1+\epsilon)(\frac{1}{n_{1}}+\frac{1}{n_{2}}).$
    Thus, $K_{1}C\leq 0$ implies
    $\frac{1}{n_{1}}+\frac{1}{n_{2}}\geq\frac{1}{1+\epsilon}>\frac{1}{1+1/3}=\frac{3}{4}.$
    So $(n_{1},n_{2})=(2,2)$ or $(2,3)$, where $(n_{1},n_{2})=(2,3)$ occurs only if $\epsilon\geq 1/5$. By the separatrix theorem, we see ${\rm Z}(\mathcal{F},C)\geq 3$.
    • If $l=0$, then $d=2$ and $d_{2}={\rm Z}(\mathcal{F},C)-1\geq 2$. Thus, $d_{2}=d=2$. This is case (1).
@@
    • If $l=1$, then $d=3$ and $d_{2}={\rm Z}(\mathcal{F},C)\geq 3$. Thus, $d=d_{2}=3$. Moreover, by the separatrix theorem, $\Xi_{1}$ must be a special $K^{>0}$-chain of type (B). Hence, we get the case (2).
  • If $l+{\rm Z}(\mathcal{F},C)-d_{2}=1$ and $k\in\{0,1\}$, then $d=k+l\in\{1,2\}$.
    • If $d=1$, then $KC=-(1+\epsilon)+{\rm Z}(\mathcal{F},C)-d_{2}\leq-\epsilon$. In this case, $C$ is the first curve of a special $K^{>0}$-chain, contradicting our assumption.
    • If $d=2$, then $k=l=1$ and ${\rm Z}(\mathcal{F},C)-d_{2}=0$. In this case, $KC=0$, and so $\Xi_{1}+C$ is a special $K^{>0}$-chain, contradicting the maximality of $\Xi_{1}$.
  • If $l={\rm Z}(\mathcal{F},C)-d_{2}=0$, then ${\rm Z}(\mathcal{F},C)=d_{2}$ and
    $k\leq d_{2}\leq d=k,$
    which imply ${\rm Z}(\mathcal{F},C)=d_{2}=k$, contradicting the separatrix theorem.
@@
In particular, cases (2-b), (2-c), (3-b) and (3-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
@@
By Lemma 4.11 and 4.12, a simple combinatorial analysis yields the desired cases, and the separatrix theorem excludes all others. ∎

### v2

@@ 4.4.2. Case II: $E$ is a $K^{=0}$-graph
If $E$ consists of a single component $C$, then $KC=0$ and $d(C)=0$. Thus $C$ must be either an irreducible elliptic Gorenstein leaf, or a smooth elliptic non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$, as described in Lemmas 4.2 and 4.3. This is case (1).
@@
By Lemma 4.2 and 4.3, each $C_{i}$ must belong to one of the following types:
@@
A direct analysis using the separatrix theorem (cf. Theorem 2.10) shows that cases (i) and (iii) cannot occur. Therefore, each $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve, and $E$ is a cycle of such curves. This is case (2). ∎
4.4.3. Case III: the general case
We now treat the remaining configurations, namely those which are neither irreducible nor $K^{=0}$-graphs.
The analysis proceeds by a successive reduction process. Starting from components with $K\cdot C<0$, we isolate certain linear subconfigurations whose numerical behaviour is completely determined. These are the special $K^{>0}$-chains classified in Proposition 4.6.
Step 1: Interaction of special chains.
If two distinct special $K^{>0}$-chains intersect, their intersection forces strong structural constraints. This situation is analyzed in Proposition 4.11, which yields an explicit finite list of admissible configurations.
@@ 4.4.3. Case III: the general case
Suppose that there exist two special $K^{>0}$-chains $F_{1}$ and $F_{2}$, whose first curves are distinct, such that they share a common component or have nonzero intersection $F_{1}\cdot F_{2}>0$. Then $E=C_{1}+\cdots+C_{r}$ is a chain belonging to one of the following cases:
@@
In particular, the corresponding point $p=\pi(E)$ is $\epsilon$-adjoint klt in all cases above.
@@
By Proposition 4.6, a finite analysis of the admissible building blocks yields the desired configurations, and the separatrix theorem (cf. Theorem 2.10) rules out any remaining possibilities.
For all components $C_{i}\subset E$, we have $KC_{i}\leq 0$, and $KC_{1}<0$. By Lemma 3.2, $E$ is a $K^{>0}$-graph, hence $p=\pi(E)$ is $\epsilon$-adjoint klt. ∎
From now on, we may therefore assume the following disjointness condition ()(*).
    $E$ is neither irreducible nor a $K^{=0}$-graph, and any two special $K^{>0}$-chains that start with different first curves are disjoint.
Step 2: Numerical reduction.
Under assumption ()(*), we can define the maximal special $K^{>0}$-chains in $E$. We adopt the following notation:
  • Let $\{\Theta_{1},\dots,\Theta_{s}\}$ denotes the set of all maximal special $K^{>0}$-chains of type (A) contained in $E$.
  • Let $\{\Xi_{1},\dots,\Xi_{t}\}$ denotes the set of all maximal special $K^{>0}$-chains of types (B), (C), or (D) contained in $E$.
We subtract the contribution of all maximal special $K^{>0}$-chains from $K$ and obtain a modified divisor
@@
We then analyze the components $C$ of $E$ not contained in any special $K^{>0}$-chains that satisfy $K_{1}\cdot C\leq 0$. We begin with the case where $C$ is not $\mathcal{F}$-invariant.
@@
Under assumption ()(*), let $C\subset E$ be an irreducible non-$\mathcal{F}$-invariant curve not contained in any special $K^{>0}$-chain. If $K_{1}C\leq 0$, then one of the following holds:
@@
    $C$ is a smooth rational curve with ${\rm tang}(\mathcal{F},C)=0$, and $E$ is an $\mathcal{F}$-star fork (cf. Definition 2.22) of type $(l;m_{1},\dots,m_{l})$ centered at $C$, where $(l;m_{1},\dots,m_{l})$ is one of the following:
@@
    In particular,
    $d(C)=d_{2}(C)=l,\qquad K_{1}C\begin{cases}<0,&\text{for case (a)},\\[2.84526pt] =0,&\text{for cases (b) and (c)}.\end{cases}$
@@
    In this case, $K_{1}C=0$ and $C$ is disjoint from all special $K^{>0}$-chains except $\Gamma_{1}$ and $\Gamma_{2}$.
@@
We first observe that $d(C)\geq 1$. If $FC>0$ for some special $K^{>0}$-chain $F$, then $F$ must be of type (C) by Proposition 4.6 and Corollary 4.8.
Without loss of generality, assume that $C$ meets transversely the chains $\Xi_{1},\dots,\Xi_{l}$. Let $m_{j}$ denote the absolute value of the determinant of the intersection matrix of $\Xi_{j}$. Since each $\Xi_{j}$ is of type (C), for $j\leq l$ we have $m_{j}\geq 2$. By (4.3), we have
@@
$\geq{\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+(1+\epsilon)d_{1}+\epsilon(d_{2}-l)+\frac{\epsilon l}{2},$
@@
Observe that if ${\rm tang}(\mathcal{F},C)\geq 1$, $p_{a}(C)\geq 1$, or $d_{1}\geq 1$, then $K_{1}C>0$ (since $d=d_{1}+d_{2}\geq 1$ and $\epsilon\in(0,1/3)$). Hence we may assume
${\rm tang}(\mathcal{F},C)=p_{a}(C)=d_{1}(C)=0,$
so that
$K_{1}C=\epsilon(d_{2}-l-2)+\sum_{j=1}^{l}\epsilon\left(1-\frac{1}{m_{j}}\right)\geq\frac{\epsilon}{2}(2(d_{2}-l)+l-4).$
Set
$x=d_{2}-l\,(\geq 0).$
Then $K_{1}C\leq 0$ gives
(4.6) $\epsilon(2x+l-4)\leq 0.$
On the other hand,
(4.7) $KC=\epsilon(x+l-2)\geq 0,$
because otherwise $C$ would start a special $K^{>0}$-chain, contradicting our assumption. Moreover, if equality holds in (4.7) and $l\geq 1$, then adjoining $C$ to a neighbouring $K^{>0}$-chain would produce a new special $K^{>0}$-chain, contradicting maximality. Hence,
(4.8) $\epsilon(x+l-2)=0\quad\Longrightarrow\quad l=0.$
We are thus reduced to a finite numerical problem for
$x,l\in\mathbb{Z}_{\geq 0},\qquad\epsilon\in(0,1/3),$
subject to (4.6)–(4.8). A straightforward inspection gives the only possible pairs:
$(x,l)=(2,0),\quad(1,2),\quad(0,3),\quad(0,4).$
Case $(2,0)$: Here $l=0$, $d=d_{2}=2$ and $K_{1}C=KC=0$. Moreover, $C$ is disjoint from all special $K^{>0}$-chains, giving case (3).
Case $(1,2)$: Here $l=2$, $d=d_{2}=3$, and
@@
and equality holds if and only if $(m_{1},m_{2})=(2,2)$. This corresponds to case (2).
Case $(0,3)$ and $(0,4)$: Here $l\in\{3,4\}$ and
$K_{1}C=-2\epsilon+\sum_{j=1}^{l}\epsilon\left(1-\frac{1}{m_{j}}\right)=\epsilon\left(-2+l-\sum_{j=1}^{l}\frac{1}{m_{j}}\right).$
@@
In this situation, $E=C+\Xi_{1}+\cdots+\Xi_{l}$ is an $\mathcal{F}$-star fork, giving case (1). ∎
We now consider the $\mathcal{F}$-invariant case.
@@
    In this case, $K_{1}C=0$ and $C$ is disjoint from all special $K^{>0}$-chains except $\Gamma_{1}$ and $\Gamma_{2}$.
@@
Moreover, we have:
  • Cases (1-b), (1-c), (2-b) and (2-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
  • $K_{1}C<0$ for cases (1-a) and (2-a).
@@
We first observe that $d(C)\geq 1$. Without loss of generality, assume that $C$ meets the chains $\Theta_{1},\dots,\Theta_{k}$ transversely at points $p_{1},\dots,p_{k}$, and the chains $\Xi_{1},\dots,\Xi_{l}$ at points $q_{1},\dots,q_{l}$. Set
$V:=\sum_{i=1}^{k}M(K,\Theta_{i})+\sum_{j=1}^{l}M(K,\Xi_{j}).$
@@
$VC=\sum_{i=1}^{k}\frac{1+\epsilon}{n_{i}}+\sum_{j=1}^{l}\frac{\epsilon}{m_{j}},$
where $n_{i}$ (resp. $m_{j}$) denotes the absolute value of the determinant of the intersection matrix of $\Theta_{i}$ (resp. $\Xi_{j}$). In particular, $n_{i}\geq 2$.
By (4.4), we have
@@
$\quad+\sum_{i=1}^{k}(1+\epsilon)\Bigl(1-\frac{1}{n_{i}}\Bigr)+\sum_{j=1}^{l}\epsilon\Bigl(1-\frac{1}{m_{j}}\Bigr)$
$\geq(1+\epsilon)(2p_{a}(C)-2+d-k-l)+l+{\rm Z}(\mathcal{F},C)-d_{2}+\frac{(1+\epsilon)k}{2}$
@@
If $p_{a}(C)\geq 1$, then $K_{1}C\leq 0$ forces $k=l=d=0$, which contradicts our assumption. Hence we may assume $p_{a}(C)=0$. In this case, the dual graph of $C+\sum_{i=1}^{k}\Theta_{i}$ forms a tree, which implies by the separatrix theorem that
(4.9) ${\rm Z}(\mathcal{F},C)\geq k+1.$
Set
$x=d-k-l\,(\geq 0),\qquad y={\rm Z}(\mathcal{F},C)-d_{2}\,(\geq 0).$
Then we have
$K_{1}C\geq(1+\epsilon)(x-2)+l+y+\frac{(1+\epsilon)k}{2}.$
Since $K_{1}C\leq 0$, this yields
(4.10) $(1+\epsilon)\Bigl(x+\frac{k}{2}-2\Bigr)+l+y\leq 0.$
On the other hand,
(4.11) $KC=(1+\epsilon)(x+k+l-2)+y\geq 0,$
because otherwise $C$ would start a special $K^{>0}$-chain, contradicting our assumption. Moreover, if equality holds in (4.11) and $k+l\geq 1$, then adjoining $C$ to a neighbouring special $K^{>0}$-chain would produce a new special $K^{>0}$-chain, contradicting maximality. Hence
(4.12) $(1+\epsilon)(x+k+l-2)+y=0\quad\Longrightarrow\quad k=l=0.$
Finally, we claim
(4.13) $x+l+y\geq 1.$
Indeed, otherwise $x=l=y=0$, which implies
$k\leq{\rm Z}(\mathcal{F},C)=d_{2}\leq d=k,$
so that ${\rm Z}(\mathcal{F},C)=k$, contradicting (4.9).
We are thus reduced to a finite numerical problem for
$x,k,l,y\in\mathbb{Z}_{\geq 0},\qquad\epsilon\in(0,1/3),$
subject to (4.10)–(4.13). A straightforward inspection gives the only possible quadruples:
$(x,k,l,y)=(2,0,0,0),\quad(1,2,0,0),\quad(0,2,0,1),\quad(0,2,1,0).$
@@
Case $(2,0,0,0)$. Here $k=l=0$, $d=2$, ${\rm Z}(\mathcal{F},C)=d_{2}\in\{1,2\}$, and $K_{1}C=KC=0$. Hence $C$ is either a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=d_{2}=1$ or a (2)(-2)-$\mathcal{F}$-curve with $d=d_{2}=2$, disjoint from all special $K^{>0}$-chains. This gives case (4) or (5).
Case $(1,2,0,0)$. Here $k=2$, $l=0$, $d=3$. We have
$K_{1}C=(1+\epsilon)\Bigl(1-\frac{1}{n_{1}}-\frac{1}{n_{2}}\Bigr),$
so that $K_{1}C\leq 0$ forces $(n_{1},n_{2})=(2,2)$, which further implies $K_{1}C=0$. Hence $C$ meets transversely two (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2. By (4.9), ${\rm Z}(\mathcal{F},C)\geq 3$, which leads to $d_{2}=d=3$. This gives case (3).
Cases $(0,2,0,1)$ and $(0,2,1,0)$. Here $k=2$, $d=l+2\in\{2,3\}$, and
$K_{1}C\geq 1-(1+\epsilon)\Bigl(\frac{1}{n_{1}}+\frac{1}{n_{2}}\Bigr),$
so that $K_{1}C\leq 0$ implies $\frac{1}{n_{1}}+\frac{1}{n_{2}}\geq 1/(1+\epsilon)>3/4$, hence $(n_{1},n_{2})=(2,2)$ or $(2,3)$, the latter only if $\epsilon\geq 1/5$. By (4.9), ${\rm Z}(\mathcal{F},C)\geq 3$.
@@
If $l=0$, then $d=d_{2}=2$ and ${\rm Z}(\mathcal{F},C)=3$, giving case (1); if $l=1$, then $d=d_{2}={\rm Z}(\mathcal{F},C)=3$, and $\Xi_{1}$ is necessarily a special $K^{>0}$-chain of type (B), giving case (2). Moreover, $K_{1}C=-\epsilon$ for case (1-a) and $K_{1}C=-\epsilon/m_{1}$ for case (2-a). ∎
Step 3: Residual configurations. By Lemmas 4.12 and 4.13, components satisfying $K_{1}\cdot C\leq 0$ are again severely restricted. If some component satisfies $K_{1}\cdot C<0$, we obtain one of finitely many explicitly described configurations. Otherwise, $E$ is a $(K_{1})^{=0}$-graph, which also reduces to a finite combinatorial classification. More precisely, we have the following proposition.
@@
Moreover, we have:
  • (i)
    cases (2-b), (2-c), (3-b) and (3-c) cannot occur if $\epsilon<\tfrac{1}{5}$.
  • (ii)
    the corresponding point $p=\pi(E)$ is $\epsilon$-adjoint klt for cases (1-a), (2-a) and (3-a).
  • (iii)
    the corresponding point $p=\pi(E)$ is not $\epsilon$-adjoint klt for cases (1-b) and (1-c).
@@
If there exists an irreducible curve $C\subset E$ as in Lemma 4.12(1) or Lemma 4.13(1)(2), then we are in cases (1)–(3).
Otherwise, no such curve exists. Hence $E$ is a $(K_{1})^{=0}$-graph. By the preceding lemmas, $E$ is obtained by combining configurations of the following types:
  • special $K^{>0}$-chains;
  • configurations as in Lemma 4.12(2)(3);
  • configurations as in Lemma 4.13(3)(4)(5).
Nevertheless, any such $E$ is ruled out by the separatrix theorem (cf. Theorem 2.10) and assumption ()(*).
Next, we prove the last part.
(i) follows from Lemma 4.13.
(ii) For cases (1-a), (2-a) and (3-a), we see $K_{1}E_{i}\leq 0$ for any component $E_{i}\subset E$ and $K_{1}C<0$ by Lemma 4.12 and 4.13. On the other hand, $E$ is connected. It follows from Lemma 3.2 that $E$ is a $(K_{1})^{>0}$-graph, so $K^{>0}$-graph. This is equivalent to saying that $p=\pi(E)$ is $\epsilon$-adjoint klt.
(iii) For cases (1-b) and (1-c), $K_{1}C=0$ by Lemma 4.12. Thus, $M(K_{1},E)=0$ and so
$M(K,E)=M(K,\Xi_{1})+M(K,\Xi_{2}).$
This implies the coefficient of $C$ in $M(K,E)$ is 0, which means $p$ is not $\epsilon$-adjoint klt. ∎